If you start at rest and accelerate to a speed of 100 miles per hour in 12 seconds, how fast will your acceleration be?
Also what steps are taken to arrive at the answer
Thanks.

Answer by mananth(16946)

If you start at rest and accelerate to a speed of 100 miles per hour in 12 seconds, how fast will your acceleration be?
...
Convert mph to ft/sec
100 mph = 100*5280/3600
= 146.66 ft/s
..
v= u+at
v = final velocity
u = initial velocit
t= time= 12 seconds
146.66= 0+a*12
divide by 12
146.66/12 = a
12.22 ft/s^2
...
m.ananth@hotmail.ca